OpenStudy (kodak.kiid):

Write 80.78 as a mixed number in simplest form.

OpenStudy (welshfella):

.78 = 78/100 can you simplify that?

OpenStudy (jdcap):

\[80.78=\frac{ 8078 }{ 100 }=\frac{ 4039 }{ 50 } =80\frac{ 39 }{ 50 }\]